OpenStudy (anonymous):

Which of the following shows a proportion? eight over eighteen equals twelve over twenty-seven three over eight equals ten over fifteen six over four equals eight over twelve twelve over twenty equals ten over eighteen

OpenStudy (texaschic101):

8/18 = 12/27 (cross multiply) (8)(27) = (18)(12) 216 = 216 (correct) -- this is a proportion 3/8 = 10/15 (cross multiply) (3)(15) = (8)(10) 45 = 80 (incorrect) -- this is not a proportion If the final result is equal, then it is a proportion. If they are not equal, then it is not a proportion.